The Pareto distribution has been used in economics as a model for a density function with a slowly decaying tail:

f(x|x0, θ) = θx0θ x−θ−1, x ≥ x0,θ > 1

Assume that x0 > 0 is given and that X1,...,Xn is an i.i.d. sample.

find a sufficient statistic for θ
Answer #1

f(z!.ro. θ)-82.02.-9-1 for x ge x_0, heta > 1

The joint pdf for X = (X1, X2,..., Xn) is,

-0-1 7l 71 71 f(z|θ) =Пе.rgri-4-1 =t?4П rì-9-1 = (θη4) | Π.r.) =g(t,0)h(z) i=1

where,

t = prod_{i=1}^{n}x_i

g (t , heta)= left ( heta^n x_0^{n heta} ight ) t^{- heta-1}

h(x) = 1

Thus, by factorization theorem,

T(X) = prod_{i=1}^{n}X_i

is a sufficient statistic for heta